OpenStudy (cenaida):

Add and select the answer in simplest form: 5/7 + 1/7 = A. 4/7 B. 5/7 C. 6/7 D. 6/14

OpenStudy (anonymous):

just this? \[5/7 = \frac{ 5 }{ 7 }\] \[1/7 = \frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

so 1/7 + 5/7 = \[\frac{ 5+1 }{ 7 } = \frac{ 6 }{ 7 } = 6/7\]

OpenStudy (anonymous):

C is correct madam :)

OpenStudy (anonymous):

yes c is ans

OpenStudy (cenaida):

Thank you! I have another question, 11/14 – 3/14 = will that equal 8/14?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

9/4! \[\frac{ 11 }{ 4 } - \frac{ 3 }{ 4 } = \frac{ 11-3 }{ 4 } = \frac{ 9 }{ 4 } = 9/4\]

OpenStudy (anonymous):

aorry 8/4! XD

OpenStudy (cenaida):

But the only options are, A. 7/14 B. 1/2 C. 4/7 D. 9/14

OpenStudy (anonymous):

8/4 is 2

OpenStudy (anonymous):

4/7 is the answer

OpenStudy (anonymous):

8/2/14/2 k

OpenStudy (cenaida):

Thanks figa! how about this question? Which of the following would you address LAST when simplifying an expression? A. Addition B. Exponents C. Multiplication D. Parentheses

OpenStudy (anonymous):

a addition

OpenStudy (cenaida):

Thank you soo muchh! and this one?? Add and write answer in simplest form: 9/11 + 1/11 = A. 8/11 B. 10/11 C. 10/22 D. 10/23

OpenStudy (cenaida):

is it B ?

OpenStudy (anonymous):

10/11

OpenStudy (anonymous):

its b@e.ali

OpenStudy (anonymous):

@cenaida dont post question within question k

OpenStudy (cenaida):

And this one? is it A ? Subtract and write answer in simplest form: 9/10 – 5/10 = A. 2/5 B. 4/10 C. 1/2 D. 3/5

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ok,wait.

OpenStudy (anonymous):

see the rules @ceinada .dont post question within question and not the complete answer

OpenStudy (anonymous):

it's b...4/10 ... @figa is right... and the last one was 10/11

OpenStudy (anonymous):

@figa is right.